# Fast LISA likelihood approximations by downsampling

### v1

@@ 3.5 Summary
The downsampling procedure we have devised, though not necessarily programmed precisely as described here for reasons of computational efficiency and/or algorithmic practicality, will consist of the following steps:
@@
    Compute the FIM of the full LISA posterior and the whitening matrix $\mathbf{C}^{(-1/2)}$ from the power spectral density (PSD)
@@
    Generate an $N_{\mathrm{f}}$-dimensional vector, $\mathbf{k}=(0,0,0,1,0,1,0,0,1,0,0,...)$, containing $N_{\mathrm{s}}$ 1’s and $(N_{\mathrm{f}}-N_{\mathrm{s}})$ 0’s, randomly positioned (or equivalently, a matrix $\mathbf{D}$ as in (4))
@@
    Use $\mathbf{k}$ to define the (whitened) $N_{\mathrm{f}}$-dimensional, downsampled residual; $\bar{\mathbf{r}}_{N_{\mathrm{f}}}=\mathbf{C}^{(-1/2)}\mathbf{r}\odot\mathbf{k}$. This seems somewhat more natural to implement for efficient computation, given the whitening required. However, an $N_{\mathrm{s}}$-dimensional version, $\bar{\mathbf{r}}_{N_{\mathrm{s}}}=\mathbf{D}\mathbf{C}^{(-1/2)}\mathbf{r}$ is perhaps conceptually clearer for visualising the downsampling
@@
    Define a new inner product on residuals/(log-)likelihood function, as in either (19) or (29). This should be approximately equal to the (log-)likelihood function defined with the full dataset for all $\bm{\theta}$, and is what shall be tested in the following section.
@@ 4 Posterior Production and Analysis
Intuitively, the fewer samples required to define a particular likelihood function, the faster likelihood evaluation and PE convergence will be; recall faster PE convergence is the sole reason for attempting downsampling. One can expect that if $N_{\mathrm{s}}\approx N_{\mathrm{f}}$, for example, dropping just a few samples from a dataset $\mathbf{d}_{\mathrm{f}}$, with, say, $\mathrm{dim}(\mathbf{d}_{\mathrm{f}})\!\approx\!10^{7}$, the downsampled posterior would be hardly distinguishable from the LISA posterior; the remaining data will still just as well constrain the likely parameters of the signal (if the noise on the remaining samples is adjusted appropriately) but the speed improvement would be slight, due to evaluating the signal model at slightly fewer points. As one drastically decreases the number of samples to only a few however, one can expect the remaining data to fit other combinations of parameters, leading to the appearance of new posterior modes; the posterior becomes more highly structured. Consider decimation by a degree such that the signal is sampled below its Nyquist frequency; aliasing should occur at this point. As mentioned, random sampling (as opposed to decimation) does very well at ameliorating this problem. However, extremely low numbers of samples from random downsampling can still cause a posterior to ‘break down’ (that is, to begin to differ qualitatively from the truth by an unacceptable degree). What we would then like to determine is: how can we quantify this breakdown of the qualitative structure of the posterior, and what is the smallest number of samples required to faithfully reproduce the true posterior, before breakdown occurs?
Changing dependencies of the posterior function (e.g. by downsampling) inevitably alters the posterior. It is not obvious how to establish criteria for how much variation may be declared ‘acceptable’; this could depend on the applications of the analysis and what a researcher hopes to gain by studying a posterior. We define our limits to be suitable for the analyses of the LISA BHB PE environment that is of interest to us. Before introducing limits in Section 4.3, we require some introduction to sources of variation in a posterior, or ‘posterior noise’, given in Section 4.1, and a measure of posterior comparison, introduced in Section 4.2.
@@ 4.1.1 Sampler noise
It generally takes an inordinate amount of effort to evaluate the posterior over a lattice, say, to any useful degree of accuracy for a realistic number of dimensions, and posterior estimators called samplers are used to explore the parameter space, building an approximate representation of the function until convergence criteria are met [19, 20, 21]. These estimates can vary to a small degree from sampler settings, between different samplers, and by the random processes that samplers use to explore the parameter space. We tested the Markov chain Monte Carlo (MCMC) sampler emcee [19] and nested samplers CPnest [20] and Nessai [22] to ensure good agreement with the main sampler we chose to provide our posterior estimates: Nessai. Using this sampler with the same settings to produce multiple posterior estimates from the same data/likelihood function, the only source of variation present results from the sampler’s internal random processes. This shall be referred to as sampler noise and is unavoidable, inducing a noise floor that cannot be overcome.
4.1.2 Downsampling datapoint selection noise
Each data sample contains different information, in the form of constraints, on the model parameters. If $N_{\mathrm{s}}$ samples are selected from an original set of $N_{\mathrm{f}}$ samples, where $N_{\mathrm{s}}\gg 1$ (and the signal model is slowly evolving) then the selection will be likely to be a good representative sample of the information present in the full dataset. However, there will of course be some variations on the parameter constraints depending on the precise selection of the $N_{\mathrm{s}}$ samples. Thus, different random datapoint choices (i.e., different realisations of $\mathbf{k}$ or $\mathbf{D}$) will incur some degree of variation of the posterior, which we shall refer to as downsampling noise. As $N_{\mathrm{s}}$ is decreased, the datapoint selection becomes less likely to contain a good representation of the information present in the original dataset. Figure 4 provides visual example of convergence of the posterior as $N_{\mathrm{s}}$ is increased.
@@ 4.1.3 Physical noise
Real data contains detector noise. Different realisations of detector noise significantly alter the posterior, the largest variation (as measured by divergence from the zero noise realisation posterior) being a shift in the position of the maximum likelihood estimate (MLE). Higher moments of the distribution contribute smaller variations.
@@ 4.2 Approximating closeness of distributions
The Kullback-Liebler (KL) Divergence is a well-known (non-symmetric) measure of distance between two distributions, defined by
@@
where $p$ and $q$ are distributions of the continuous variable $\bm{\theta}\in\Theta$, where $\Theta$ is the parameter space. Here, $p$ is thought of a the target/observed/correct distribution, and $q$ the hypothesis distribution. The Jensen-Shannon (JS) Divergence, defined as
@@
is a symmetric distance measure (and a metric) between distributions. It will be better-suited for a subset of our requirements, where the strict target distribution is unattainable.
As mentioned, the posterior distributions themselves are unfortunately computationally very expensive. For high dimensional parameter spaces, the divergences are vastly more expensive still; first, the posteriors are required to be known, then one needs to perform the further mathematical operations as per the definitions. We must resort to approximation techniques.
@@ 4.2.1 Combined distances of marginal distributions
Consider using the divergences of marginalised distributions. Marginalisation is certainly not ideal; as it is a projection, information is lost. However it can still be useful as a measure to show when variations between the marginals of posteriors become very small, and thus point to convergence, and divergences between 1-dimensional distributions are inexpensive to compute. One has an infinite choice of ways in which to project a distribution with $\geq 2$ dimensions into 1-dimension (by parameter mixing). We are usually interested in the unmixed model parameter constraints however, so it is sensible to choose the projections that are the marginalisations onto the usual, unmixed model parameters.
Since the JS divergence can be used to define a metric and thus a distance, we can combine distances of these marginalisations to form a new distance. We can then define a metric between distributions $p,q$ as:
$d_{\mathrm{JS}}(p,q)\equiv\sqrt{D_{\mathrm{JS}}(p\,||\,q)}\,.$ (32)
Let us now define a generalised Pythagorean distance function
$\overline{d}_{\mathrm{JS}}(p,q)\equiv\sqrt{\frac{1}{k}\sum_{i=0}^{k-1}d^{2}_{\mathrm{JS}}\left(O^{i}(p),O^{i}(q)\right)}$ (33)
@@
$O^{i}(p)\equiv\int_{\hat{\Theta}_{i}}p(\bm{\theta})\;\mathrm{d}\theta^{1}...\mathrm{d}\theta^{i-1}\mathrm{d}\theta^{i+1}...\mathrm{d}\theta^{k}\,,$ (34)
projects the distribution $p$ onto the $i^{\mathrm{th}}$ parameter axis, where $\hat{\Theta}_{i}\cong\mathbb{R}^{k-1}$ is the space spanned by vectors $\hat{\bm{\theta}}_{i}=(\theta^{1},...,\theta^{i-1},\theta^{i+1},...,\theta^{k})$. The function $\overline{d}_{\mathrm{JS}}$ satisfies the requirements of a metric and is thus a metric. If we square the new metric to define a new divergence $\overline{D}_{\mathrm{JS}}$, we will have that
$\overline{D}_{\mathrm{JS}}(p\,||\,q)\equiv\frac{1}{k}\sum_{i=0}^{k-1}D_{\mathrm{JS}}\left(O^{i}(p)\,||\,O^{i}(q)\right)\,.$ (35)
The range of $D_{\mathrm{JS}}$ is $[0,1]$, the factor of $1/k$ introduced above thus ensures that the range of $\overline{D}_{\mathrm{JS}}$ is $[0,1]$ also. We shall refer to this divergence, which will be instrumental for indicating the convergence of downsampled posteriors, as the combined marginal Jenson-Shannon (CMJS) divergence. We will similarly define the combined marginal Kullback-Liebler (CMKL) divergence as
$\overline{D}_{\mathrm{KL}}(p\,||\,q)\equiv\frac{1}{k}\sum_{i=0}^{k-1}D_{\mathrm{KL}}\left(O^{i}(p)\,||\,O^{i}(q)\right)\,.$ (36)
@@ 4.3 Acceptance criteria
We are almost ready to propose criteria with which to determine the minimum number of samples required to define an acceptable downsampled posterior. However, there is no preexisting benchmark, and some researcher may require more or less accuracy than another. We lay out a particular choice, suitable for recreating the LISA BHB PE test environment up to an accuracy of posterior morphology defined by physical noise, which is in line with our purpose for this work. Recall that we have three types of posterior noise to consider: physical noise, sampler noise and downsampling noise. We first determine how to quantify noise between distributions in Section 4.3.1, use the accuracy of the sampler to "bootstrap" up to an acceptable approximate target posterior in Section 4.3.2, then use physical noise to define the accuracy cutoff for downsampled posteriors in Section 4.3.3.
@@ 4.3.1 Induced distribution noise
Using the approximate divergences outlined in the previous section, we can begin to characterise some basic statistical properties of the variation in the closeness of distributions, arising from some particular noisy process (i.e. the posterior noise sources). We now define two one-parameter (parameterised by $N_{\mathrm{s}}$) families of sets of divergences:
@@
    When a target distribution is not known, JS divergence is useful as a notion of relative distance. By generating $K$ posterior estimations, we can form a population of $\binom{K}{2}$ pairs (‘$K$ choose 2’) of posteriors, and thus a set of $\binom{K}{2}$ divergences. The set of divergences is given by
    $X_{\mathrm{est}}=\bigcup_{i<j\leq K-1}\overline{D}_{\mathrm{JS}}\left(p_{\mathrm{est}}^{i}\,\Big{|}\Big{|}\,p_{\mathrm{est}}^{j}\right)\,,$ (37)
    where $p_{\mathrm{est}}^{i}$ is the $i^{\mathrm{th}}$ element of the $K$ estimate posteriors.
@@
    When a target distribution is known, we have the notion of absolute distances where the target acts in a sense as the ‘origin’ of the KL divergence space of distributions. With $K$ posterior estimations and a target distribution $q$, we have the set of $K$ divergences given by
    $Y^{q}_{\mathrm{est}}=\bigcup_{i\leq K-1}\overline{D}_{\mathrm{KL}}\left(q\,\Big{|}\Big{|}\,p_{\mathrm{est}}^{i}\right)\,,$ (38)
    where again $p_{\mathrm{est}}^{i}$ is the $i^{\mathrm{th}}$ element of the $K$ posteriors.
Let the mean of a set of divergences, $X$, be $\mu[X]$, and the standard deviation $\sigma[X]$. Whilst indeed a Poisson distribution is better suited for representing these sets of divergences since they are always strictly positive, we can still use both statistics to quantify distribution noise. Using both mean and standard deviation, we can define more consistent constraints on the acceptability of downsampled posteriors.
4.3.2 "Bootstrapping"
Unfortunately, using the fully sampled posteriors as target distributions to compare the downsampled posteriors to is not an available option, due to the immense computational costs. We therefore require a so-called "bootstrapping" procedure, to show that our approximations are sufficiently accurate that convergence criteria defined using them remain valid. Note that we require the reasonable assumption that the posterior estimates provided by the sampler are unbiased, and that this particular bootstrapping procedure is based on the knowledge that the sampler is well-tested, and highly accurate and consistent [22]. However, this procedure is only possible as the non-zero sampler noise provides a noise baseline for comparison of the downsampling noise; the target distribution we obtain will be roughly as good a representation of the true posterior as the estimate given by the sampler. For a ‘perfect’ sampler with zero noise, no downsampling noise comparison can be made, and at the other extreme, a poorly performing sampler with high posterior noise does not allow for a good grasp of downsampling noise. In those cases some other bootstrapping procedure would be required. A description of the procedure is as follows (this bootstrapping procedure is independent of the downsampling scheme and method, any may be used for this step if they satisfies the required conditions).

### v2

@@ 3.5 Summary
The downsampling procedure we have devised, though not necessarily implemented exactly as described here (for reasons of computational efficiency and algorithmic practicality), consists of the following steps:
@@
    Compute the FIM for the full LISA posterior and the whitening matrix $\mathbf{C}^{(-1/2)}$ from the power spectral density (PSD).
@@
    Generate an $N_{\mathrm{f}}$-dimensional sample-selection vector $\mathbf{k}=(0,0,0,1,0,1,0,0,1,0,0,\dots)$ containing $N_{\mathrm{s}}$ ones and $(N_{\mathrm{f}}-N_{\mathrm{s}})$ zeros, randomly positioned (or equivalently, construct a matrix $\mathbf{D}$ as in (4)).
@@
    Use $\mathbf{k}$ to define the (whitened) downsampled residual,
    $\bar{\mathbf{r}}=\mathbf{C}^{(-1/2)}\mathbf{r}\odot\mathbf{k}\,.$
    This representation is convenient for efficient computation given the whitening step. An $N_{\mathrm{s}}$-dimensional version, $\bar{\mathbf{r}}_{N_{\mathrm{s}}}=\mathbf{D}\mathbf{C}^{(-1/2)}\mathbf{r}$, is conceptually equivalent and may be clearer for visualising the downsampling.
@@
    Define a new inner product on residuals and hence a new (log-)likelihood, as in either (18) or (69). This is constructed to be approximately equal to the (log-)likelihood defined with the full dataset for all $\bm{\theta}$, and its performance will be tested in the following section.
@@ 4 Posterior Production and Analysis
Intuitively, fewer samples mean cheaper likelihood evaluations and faster PE convergence; this is the main motivation for downsampling. If $N_{\mathrm{s}}\approx N_{\mathrm{f}}$ (e.g. dropping only a small fraction of samples from a dataset $\mathbf{d}_{\mathrm{f}}$ with $\dim(\mathbf{d}_{\mathrm{f}})\approx 10^{7}$), the downsampled posterior will be essentially indistinguishable from the full LISA posterior (provided the retained-sample noise is adjusted appropriately), but the speed-up is modest. As $N_{\mathrm{s}}$ is reduced more aggressively, the remaining data become consistent with qualitatively different parameter combinations, producing extra modes and more highly structured posteriors. For example, uniform decimation below the Nyquist rate leads to aliasing. Random sampling, rather than regular decimation, largely mitigates this, but at very low $N_{\mathrm{s}}$ the posterior can still ‘break down’ in the sense of differing qualitatively and unacceptably from the full-data posterior. Our goal is therefore to quantify the onset of this breakdown and to determine the smallest $N_{\mathrm{s}}$ that still reproduces the true posterior faithfully.
Any change to the data or likelihood definition (such as downsampling) inevitably alters the posterior, and it is not a priori clear when such changes should be regarded as ‘acceptable’, since this depends on the application and on which features of the posterior are scientifically relevant. We therefore tailor our acceptance criteria to the LISA BHB PE environment of interest here. Before introducing these criteria in Section 4.3, we first describe sources of variation, or ‘posterior noise’, in Section 4.1, and then introduce a quantitative measure of posterior distance in Section 4.2.
@@ 4.1.1 Sampler noise
It is generally infeasible to evaluate the posterior on a dense grid in parameter space for realistic dimensionalities, so we rely on stochastic samplers to construct approximate posterior estimates [emcee, CPnest, Dynesty]. These estimates can vary slightly between different samplers, between different settings of a given sampler, and between different realisations of the underlying random processes. We compared the Markov chain Monte Carlo (MCMC) sampler emcee [emcee] and the nested samplers CPnest [CPnest] and Nessai [Nessai], and found good agreement with our chosen default, Nessai. With fixed settings, repeated Nessai runs on the same data and likelihood differ only through the sampler’s internal randomness. We refer to this irreducible variation as sampler noise; it induces a noise floor that cannot be overcome.
4.1.2 Downsampling noise
Each data sample carries different information about the model parameters. If $N_{\mathrm{s}}$ samples are selected from an original set of $N_{\mathrm{f}}$ samples with $N_{\mathrm{s}}\gg 1$ (and the signal model is slowly evolving), the selected subset is likely to be a good representative of the information in the full dataset. Nonetheless, the geometry will vary slightly with the precise choice of the $N_{\mathrm{s}}$ samples. Thus different random data point selections (i.e. different realisations of $\mathbf{k}$ or $\mathbf{D}$) induce some variation in the posterior, which we refer to as downsampling noise. As $N_{\mathrm{s}}$ is decreased, a typical selection becomes less likely to provide a good representation of the information present in the original dataset.
@@ 4.1.3 Physical noise
Real data contain detector noise. Different realisations of this noise can significantly alter the posterior; the dominant effect (as measured by divergence from the zero-noise posterior) is typically a shift in the position of the maximum likelihood estimate (MLE), while changes in higher moments contribute smaller variations.
To quantify the intrinsic level of posterior-to-posterior variation induced by physical noise, we consider an ensemble of posteriors obtained from multiple independent noise realisations for the same injected signal and likelihood. For each realisation we mean-zero the posterior in parameter space. This removes the overall shifts and focuses on changes in the posterior shape. We then compute the distribution of distances between these mean-zeroed posteriors using the divergence measures introduced below. This defines a “physical-noise floor”: the typical amount of variation in posterior geometry that arises purely from changing the detector noise realisation.
The resulting noise floor is system- and SNR-dependent, and is therefore tied to the specific LISA BHB configurations studied here. Further, the precise value depends on the chosen alignment prescription (here, mean-zeroing in each parameter); this choice is fixed throughout. Within these conventions, we treat downsampling-induced distortions as acceptable whenever the corresponding posterior distances are comparable to, or smaller than, this physical-noise floor.
@@ 4.2 Approximating closeness of distributions
The Kullback-Leibler (KL) divergence is a well-known (non-symmetric) measure of discrepancy between two distributions, defined by
@@
where $p$ and $q$ are distributions of the continuous variable $\bm{\theta}\in\Theta$, with $\Theta$ the parameter space. Typically, $p$ is viewed as the target/“true” distribution, and $q$ as an approximate or hypothesised distribution.
The Jensen-Shannon (JS) divergence is defined as
@@
and is a symmetric, bounded divergence between distributions. Moreover, its square root, $\sqrt{D_{\mathrm{JS}}(p||q)}$, defines a metric on the space of distributions. This will be better suited for our requirements, particularly in situations where the exact target distribution is unattainable.
As mentioned, the posterior distributions themselves are computationally very expensive to evaluate accurately, and for high-dimensional parameter spaces the exact divergences are more expensive still: one must first estimate the posteriors, and then perform the additional integrals implied by the definitions above. We therefore resort to approximation techniques in order to obtain practical measures of the closeness between posteriors.
@@ 4.2.1 Combined distances of marginal distributions
Consider using divergences between marginal distributions. Marginalisation is not ideal; being a projection, it necessarily discards information, but it remains useful as a low-cost diagnostic: if the 1D marginals of two posteriors become extremely similar, this strongly suggests convergence, and divergences between 1D distributions are relatively inexpensive to compute. There are infinitely many ways in which to project a distribution with $\geq 2$ dimensions into 1-dimension (by parameter mixing). We are usually interested in the unmixed model parameter constraints however, so it is natural to choose the projections that are the marginalisations onto the unmixed model parameters.
The Jensen-Shannon divergence $D_{\mathrm{JS}}(p\|q)$ induces a metric
$d_{\mathrm{JS}}(p,q)\equiv\sqrt{D_{\mathrm{JS}}(p\|q)}$
on the space of 1D distributions. We now combine these marginal distances into a single quantity by taking a weighted Euclidean norm:
$\overline{d}_{\mathrm{JS}}(p,q)\equiv\sqrt{\sum_{i=0}^{k-1}W_{i}\,d_{\mathrm{JS}}^{2}\!\left(O^{i}(p),O^{i}(q)\right)},$ (21)
@@
$O^{i}(p)\equiv\int_{\hat{\Theta}_{i}}p(\bm{\theta})\,\mathrm{d}\theta^{1}\cdots\mathrm{d}\theta^{i-1}\mathrm{d}\theta^{i+1}\cdots\mathrm{d}\theta^{k}$ (22)
projects the distribution $p$ onto the $i$-th parameter axis, and $\hat{\Theta}_{i}\cong\mathbb{R}^{k-1}$ is the space spanned by $\hat{\bm{\theta}}_{i}=(\theta^{1},\ldots,\theta^{i-1},\theta^{i+1},\ldots,\theta^{k})$.
We define the weights
$2W_{i}\equiv\frac{\mathcal{H}_{O^{i}(p)}}{\sum_{j=0}^{k-1}\mathcal{H}_{O^{j}(p)}}+\frac{\mathcal{H}_{O^{i}(q)}}{\sum_{j=0}^{k-1}\mathcal{H}_{O^{j}(q)}}\,$, (23)
where $\mathcal{H}_{\mathcal{X}}$ denotes the (Shannon) entropy of the distribution $\mathcal{X}$. Thus each marginal contributes to the combined distance according to its relative contribution to the total marginal entropy of $p$ and $q$. With this choice we have $\sum_{i}W_{i}=1$.
Since $d_{\mathrm{JS}}$ is a metric on each marginal and $\overline{d}_{\mathrm{JS}}$ is a weighted $\ell^{2}$ combination of these marginal distances, it defines a metric on the space of $k$-tuples of 1D marginals. When regarded as a function of the full joint posteriors $p$ and $q$, it should formally be viewed as a pseudometric, because distinct joint distributions may share the same set of marginals. This limitation is acceptable for our purposes, as we use $\overline{d}_{\mathrm{JS}}$ purely as a convergence diagnostic based on parameter-wise constraints.
Squaring the combined distance yields a corresponding divergence
$\overline{D}_{\mathrm{JS}}(p\|q)\equiv\sum_{i=0}^{k-1}W_{i}\,D_{\mathrm{JS}}\!\left(O^{i}(p)\|O^{i}(q)\right),$ (24)
which we refer to as the combined marginal JS (CMJS) divergence. With logarithms taken in base 2, each $D_{\mathrm{JS}}$ lies in $[0,1]$, and the normalisation $\sum_{i}W_{i}=1$ ensures that $\overline{D}_{\mathrm{JS}}\in[0,1]$ as well. Analogously, we define the combined marginal Kullback-Leibler (CMKL) divergence as
$\overline{D}_{\mathrm{KL}}(p\|q)\equiv\sum_{i=0}^{k-1}W_{i}\,D_{\mathrm{KL}}\!\left(O^{i}(p)\|O^{i}(q)\right).$ (25)
These combined marginal divergences will be used in Sec. 4.3 as practical measures of the difference in posterior geometry under changes in downsampling rate.
@@ 4.3 Acceptance criteria
We now propose criteria for determining the minimum number of samples required to define an acceptable downsampled posterior. There is no pre-existing benchmark, and different applications may tolerate different levels of approximation. Here we lay out a particular choice, suitable for recreating the LISA BHB PE test environment up to an accuracy in posterior geometry set by physical noise, which is in line with our purpose for this work. Recall that we have three types of posterior noise to consider: physical noise, sampler noise, and downsampling noise. We first define how to quantify noise between distributions in Section 4.3.1, use posterior variability under the sampler and downsampling to “bootstrap” an acceptable approximate target posterior in Section 4.3.2, and then use physical noise to set an accuracy threshold in Section 4.3.3.
@@ 4.3.1 Induced distribution noise
Using the approximate divergences outlined in the previous section, we can begin to characterise basic statistical properties of the variation in the closeness of distributions arising from a given noisy process (i.e. the different sources of posterior noise). We now define two one-parameter (in $N_{\mathrm{s}}$) families of sets of divergences:
@@
    When a target distribution is not known, the JS divergence is useful as a notion of relative distance. Given $K$ posterior estimates, we can form a population of $\binom{K}{2}$ pairs of posteriors and thus a set of $\binom{K}{2}$ divergences. We define
    $X_{\mathrm{est}}(N_{\mathrm{s}})=\left\{\overline{D}_{\mathrm{JS}}\!\left(p_{\mathrm{est}}^{(i)}\,\Big|\Big|\,p_{\mathrm{est}}^{(j)}\right):1\leq i<j\leq K\right\},$ (26)
    where $p_{\mathrm{est}}^{(i)}$ is the $i^{\mathrm{th}}$ estimated posterior.
@@
    When a target distribution is known, we have a notion of absolute distance in which the target plays the role of an “origin” in divergence space. With $K$ posterior estimates and a target distribution $q$, we define the set of $K$ divergences
    $Y^{q}_{\mathrm{est}}(N_{\mathrm{s}})=\left\{\overline{D}_{\mathrm{KL}}\!\left(q\,\Big|\Big|\,p_{\mathrm{est}}^{(i)}\right):1\leq i\leq K\right\}.$ (27)
Let the mean of a set of divergences $X$ be denoted $\mu[X]$, and the standard deviation by $\sigma[X]$. The empirical distributions of these divergences are strictly non-negative and typically skewed, so a Gaussian approximation is not exact; nevertheless, $\mu[X]$ and $\sigma[X]$ provide convenient summary statistics with which to quantify “distribution noise” and to formulate consistent acceptance criteria for downsampled posteriors.
4.3.2 Bootstrapping
Using fully sampled posteriors as target distributions for assessing the accuracy of downsampled posteriors is not feasible in realistic applications, because obtaining the former is prohibitively expensive. We therefore introduce a bootstrapping procedure to demonstrate that our approximations are sufficiently accurate and to define a minimal reliable downsampling rate.
The basic idea is to regard the variability of posteriors under downsampling as a function of $N_{\mathrm{s}}$, and to compare this variability against a prescribed accuracy threshold $\tau$ that encodes what we deem to be an acceptable distortion of the posterior. In Sec. 4.3.3 we instantiate $\tau$ using the mean-zeroed variability induced by different physical noise realisations, but the bootstrapping logic itself is independent of that particular choice. It only assumes that we have (i) an approximately unbiased sampler for the posterior, and (ii) a way of quantifying distances between posteriors via the divergences $Y^{q}_{\mathrm{est}}(\cdot)$ defined in Eq. (27).
Operationally, we assume that the additional downsampling noise is monotone in $N_{\mathrm{s}}$: as $N_{\mathrm{s}}$ is reduced, the downsampling-induced variability in the posterior increases (or at least does not decrease). The goal is then to identify a minimal value $N_{\mathrm{s}}^{\min}$ above which the variability induced by downsampling is always below the accuracy threshold $\tau$, and beyond which the posterior is effectively independent of further increases in $N_{\mathrm{s}}$.